The SuperB factory

physics prospects and project status

Marcin Chrząszcz

Institute of Nuclear Physics, Polish Academy of Science,

on behalf of the SuperB collaboration

21^{st} September 2012 Introduction SuperB Infrastructure

Accelerator Luminosity

B Physics Precision Measurements Rare B Physics TDCP B → X_{s}γ B Rare Decays B_{s} Decays Charm Physics

τ Physics LFV τg − 2 EDM at SuperB EDM at SuperB CP Violation

B factories

B factories have contributed to many important physics discoveries over the last decade. They will be succeeded the Super Flavor Factories:

Super Flavor Factories

1 Data 75ab^{−1} 2 Luminosity 10^{36}cm^{−2}s^{−1} 3 Flexibility to run on charm threshold with luminosity 10^{35}cm^{−2}s^{−1} 4 Longitudinal polarization of electron beam 80% 5 Upgraded BaBar detector 6 Start of data taking: 2018 7 10ab^{−1} per year

The SuperB factory  Introduction  3 / 27

Tor Vegata Site

Important dates:

1 TDR: Autumn this year. 2 Colliding beams: June 2018.

Tor Vegata Site Quest for Luminosity

L ∝ \frac{1}{\sqrt{β}_{y}} , Φ ≈ \frac{σ_{z}}{σ_{x}} \frac{θ}{2}

CKM Matrix

{Δ\overline{η} = 0.016 \atop Δ\overline{ρ} = 0.028}

CKM Matrix

Δ\overline{η} = 0.016 Δ\overline{ρ} = 0.028 Δ\overline{η} = 0.0024 Δ|V_{cb}|_{incl} = 0.5% Δ|V_{cb}|_{excl} = 1.0% Δ\overline{ρ} = 0.0028 Δ|V_{ub}|_{incl} = 1.0% Δ|V_{ub}|_{excl} = 3.0%

B → τν

Precise SM prediction:

Br(B → lν) = \frac{G_{F}^{2}m_{B}}{8π}m_{l}^{2}(1 − \frac{m_{l}^{2}}{m_{B}^{2}})f_{B}^{2}|V_{ub}|^{2}τ_{B} In SUSY: Br(B → lν) = \frac{G_{F}^{2}m_{B}}{8π}m_{l}^{2}(1 − \frac{m_{l}^{2}}{m_{B}^{2}})f_{B}^{2}|V_{ub}|^{2}τ_{B}(1 − \frac{tan^{2}β}{1 + \overline{ϵ}tanβ} \frac{m_{B}^{2}}{m_{H}^{2}} )

Time-Dependent CP (TDCP)

Time-dependent CP analysis can show signs of new physics. One has to study a set of modes: b → s\overline{s}c, b → s Current experimental results show Δ(SM - Observed): Δsin(2β) = 2.7σ, penguin Δsin(2β) = 2.1σ, tree Golden modes in SuperB: B → J/ψK^{0}, B → η′K^{0}, B → f_{0}K_{s}^{0}

B → X_{s}γ

Very important probe for new physics! Current experimental average: Br(B → X_{s}γ) = (3.52 ± 0.23 ± 0.09)10^{−4} Theoretical prediction from NNLO: Br(B → X_{s}γ) = (3.15 ± 0.23)10^{−4} There are two ways to study this decay:

1 Exclusive:

• The earliest results were done using a large number of exclusive decays, which were fully reconstructed

• Errors arising from unseen modes

• Obsolete for SuperB

2 Inclusive:

• Use tagging to tag the other B

• No requirements on X_{s}

• Disadvantage: Cut on photon energy

• Effort to keep the cut as small as possible

Experimentally challenging to measure inclusive decays.

The SuperB factory  B Physics  10 / 27

B Rare Decays

B^{±} → D^{(∗)}τ^{±}ν Babar ref. arXiv:1205.5442 Hot decay for SuperB!

Observables:

• R(D) = \frac{B → Dτν}{B → Dℓν}

• R(D^{∗}) = \frac{B → D^{∗}τν}{B → D^{∗}ℓν}

R(D)  R(D^{∗})

BaBar  0.440 ± 0.071 0.332 ± 0.029 SM  0.297 ± 0.017 0.252 ± 0.003 Difference 2.0σ  2.7σ The SuperB factory  B Physics  11 / 27

B_{s} Decays

B_{s} is clearly LHCb domain

Short runs at CLEO and Belle showed that e^{+} e^{−} can also contribute in B_{s} studies

Observable  Error on 1fb^{−1} Error on 30fb^{−1}
ΔΓ[ps^{−1}] 0.16 0.03
β_{s} from B_{s} → J/ψϕ[deg] 16  6
β_{s} from B_{s} → K\overline{K}^{0}[deg] 24  11
| \frac{V_{td}}{V_{ts}} |  0.08  0.017

Potentials in SuperB:

1 Decays with neutral particle B_{s} → J/ψη , B_{s} → K_{S}^{0}π, B_{s} → D^{∗}K_{S}^{0}, B_{s} → Φη′ 2 Measurements of B(B → γγ). SM prediction

B(B → γγ) = (2 − 4) × 10^{−7}. NP (SUSY) B(B → γγ) = 5 × 10^{−6}.

3 Measurements of semi-leptonic asymmetry. A_{SL}^{s} = \frac{1 − |\frac{q}{p} |^{4}}{1 + |\frac{q}{p} |^{4}} = \frac{N_{1} − N_{2}}{N_{1} + N_{2}}

N_{1} = B_{s} → \overline{B}_{s} → D_{s}^{∗−}ℓ^{+}ν N_{2} = B_{s} → \overline{B}_{s} → \overline{D}_{s}^{∗}ℓ^{−}ν

Charm Physics

1 Plan for running at ψ(3770) threshold 2 Scenario: Collect 500fb^{−1} 3 D tag possible; other meson can be studied with very small

background Potential improvement from SuperB:

• Improved measurement of the mixing parameters x_{D} and y_{D}

• CP violation in \overline{D} − \overline{D}: A_{SL} = \frac{N_{1} − N_{2}}{N_{1} + N_{2}} N_{1} = Γ(D^{0} → ℓ^{−}νK^{+}), N_{2} = Γ(\overline{D^{0}} → ℓ^{+}νK^{−})

• Search for D^{0} → μμ

• Quantum correlations allow one to measure relatively strong phase

The SuperB factory  B Physics  13 / 27

τ Physics

1 Lepton Flavaur Violation

• Some SUSY model can occure in SuperB sensitive.

• Complementary to searches in LHC and MEG.

• Golden channels: τ → 3ℓ, τ → ℓγ, τ → ρℓ, τ → ℓη .

2 τg − 2

• MSSM can explain 3 × 10^{−9} discrepancy.

• SuperB sensitivety is in range of doing this.

3 τ EDM and CPV

• In SuperB sensitivity!

• τ EDM constrained by electron EDM upper limit to a range inaccessible by SuperB

CMSSM Model

• N_{i} - right handed neutrinos • ν_{i} - left handed neutrinos • ϕ_{i} - complex mixing angle • ϕ_{13} - PNMS matrix.

• LFV up to present limit

• τ → μγ complementary to μ → eγ

JHEP11(2006)090

The SuperB factory  τ Physics  15 / 27

NUHM Model

• δ_{1}, δ_{2} parametrizes the non universal Higgs masses.

arXiv:0812.2692v1

• Increase sensitivity for τ → f_{0}(980)μ, τ → ημ, than to τ → μγ

JHEP11(2006)090

The SuperB factory  τ Physics  16 / 27

SuperB sensitivity

1 Taking the BaBar analysis results and improving:

\sqrt{L_{SuperB}/mathcalL_{BaBar}} ≈ 12 2 Signal is risening linearly: L_{SuperB}/mathcalL_{BaBar} 3 Sensitivety increases with detector resolution. 4 Babar papers used to extrapolate:

• Phys.Rev.Lett.104:021802,2010, arXiv:0908.2381v2

• PhysRevD.81.111101(2010), arXiv:1002.4550v1

τ → ℓγ Sensitivity

• Better tracking resolution, reduced Δm − ΔE box by 65%

• Higher photon efficiency

• Increase of geometry acceptance

• Thicker signal peak

• Approximate frequentistic upper limits, only Poissonian BKG uncertainty.

• Smaller boost improves the performance of the fit.

SuperB limits: Process Error on 90% upper limit 3σ observation τ → μγ 2.4 × 10^{−9}  5.4 × 10^{−9} τ → eγ 3.0 × 10^{−9}  6.8 × 10^{−9}

Polarization

1 SuperB will have polarized

electron beam (80%) 2 One can use this information in

NP searches 3 TAUOLA SUSY decay model. 4 Discriminating between NP

models!

SuperB sensitivity for τ → 3ℓ

1 Taking the BaBar analysis results and improving:

\sqrt{L_{SuperB}//mathcalL_{BaBar}} ≈ 12 2 Signal is risening linearly: L_{SuperB}/mathcalL_{BaBar} 3 No detector resolution assumed. 4 Approximate frequentistic upper limits, only Poissonian BKG

uncertainty 5 Babar papers used to extrapolate:

• Phys.Rev.Lett.104:021802,2010, arXiv:0908.2381v2

• PhysRevD.81.111101(2010), arXiv:1002.4550v1

τ → 3ℓ

τ → eee  τ → e^{−}e^{+}μ

τ → e^{−}e^{−}μ

τ → 3ℓ

τ → μμμ  τ → μ^{−}μ^{+}e

τ → μ^{−}μ^{−}e

LFV Summary

Current analysis:

• SuperB will be the cutting age factory for LFV in τ decays.

• Beam polarization will improve the the analysis and make distinguishng among NP models possible.

Process Error on 90% upper limit 3σ observation
τ → μγ 2.4 × 10^{−9}  5.4 × 10^{−9}
τ → eγ 3.0 × 10^{−9}  6.8 × 10^{−9}

τg − 2

• MSSM would shift muon g − 2 by about the presently observed discrepancy Δa_{μ} ≈ 3 × 10^{−9}.

• SuperB sensitivity estimates: σ(a_{τ} = 2.6 × 10^{−6}.

JHEP098P1108

• SuperB measures a_{τ} (q^{2}) from final state distributions of e^{+}e^{−} → τ^{+}τ^{−} See M.Passera talk

• Luckly NP contributions are constant for small q^{2}.

The SuperB factory  τ Physics  24 / 27

EDM at SuperB

• Experimental status: |d_{e}| < 1.6 × 10^{−}27.

PhysRevLett.88.071805

• NP expect: |d_{τ} | ∝ (m_{τ} /m_{e})|d_{e}|

• SuperB upper limit |d_{e}| ≈ 10^{−22} SuperB 2010 Physic Report

• Again we measure |d_{e}|(q^{2}).

• Luckly NP contributions are constant for small q^{2}.

EDM at SuperB

Belle result:

1 29.5fb^{−1} data sample 2 Resolution: 0.9 − 1.7 × 10^{−19}ecm

3 J. Bernabeu hep-ex/0210066

4 Extrapolation for SuperB (75ab^{−1}): σ(d_{τ}) = 17 − 34 × 10^{−17}ecm. 5 No beam polarization assumed!

Other aproach: arXiv:0707.1658v1 • Assume beam polarity: (80 ± 1).

• 80% geometry acceptance.

• Track reconstrucion 97.5%.

• σ(d_{τ}) ≈ 10 × 10^{−17}ecm

CP Violation

• CP violation has never been observed in τ sector

• SM prediction is negligibly small O(10^{−12})l in τ^{±} → K^{pm}π^{0}ν.

• Any observation is clear indication of NP

• Very few NP models can explain this:

1 RPV SUSY

2 Multi Higgs models

• SuperB can improve sensitivity 75 times compared to CLEO

Backup

Backup

The SuperB factory  τ Physics  28 / 27